OpenStudy (anonymous):

Determine the sixth roots of 8 as well as of -8

OpenStudy (lukecrayonz):

for 8, sqrt(2), for -8, -sqrt(2)

OpenStudy (anonymous):

you want the six complex roots right?

OpenStudy (anonymous):

yes the six complex roots

OpenStudy (anonymous):

lets do it the snap way draw the unit circle. then we can easily find the sixth roots of 1 once we have those, we multiply each by \[\sqrt{2}\] because \[\sqrt{2}^6=2^3=8\]

OpenStudy (anonymous):

since we know that \[1^6=1\] we know one sixth root of 1, namely 1. divide the unit circle into six equal parts, where one is at (1,0)

OpenStudy (anonymous):



OpenStudy (anonymous):

working in degrees you can say 360/6=60 or in radians \[\frac{2\pi}{6}=\frac{\pi}{3}\] so the coordinates of each of these points should be easy to find. first one is \[\cos(60)+i\sin(60)=\frac{1}{2}+\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

others also easy to find \[-\frac{1}{2}+\frac{\sqrt{3}}{2}i,-1,-\frac{1}{2}+-\frac{\sqrt{3}}{2}i,\frac{1}{2}-\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

multiply each of these by \[\sqrt{2}\] and you are done

OpenStudy (lukecrayonz):

Oh, i thought he just wanted 8^(1/6) hahaha